# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 2.2. Affine transformations
We will also need the following alternative interpretation of $\operatorname{Aff}(M)$. Let $\widetilde{M}=M\oplus\mathbb{Z}$ with projection map $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ onto the last coordinate. Then $M\times\{1\}=\operatorname{pr}^{-1}(1)$ is an affine lattice. Moreover, if $0_{M}$ denotes the origin in $M$, then we may use the distinguished point $(0_{M},1)$ in $M\times\{1\}$ to identify $\operatorname{Aff}(M)$ with $\operatorname{Aff}(M\times\{1\})$. On the other hand, $\operatorname{Aff}(M\times\{1\})$ is isomorphic (via restriction) to the subgroup $\{\widehat{\phi}\in\operatorname{GL}(\widetilde{M}):\operatorname{pr}\circ\widehat{\phi}=\operatorname{pr}\}\subset\operatorname{GL}(\widetilde{M})$ i.e. the subgroup of $\operatorname{GL}(\widetilde{M})$ that preserves the affine lattice $M\times\{1\}$. In this way, we will often view $\operatorname{Aff}(M)$ as a subgroup of $\operatorname{GL}(\widetilde{M})$. Explicitly, if $\nu(u)=A(u)+b$ for some $A\in\operatorname{GL}(M)$, $b\in M$ and any $u\in M$, then the corresponding linear transformation is $\widehat{\phi}(u,m)=(A(u)+mb,m)$ for any $u\in M$ and $m\in\mathbb{Z}$.
@@ 2.3. Representation theory of finite groups
By Lemma 2.5, there is a $g$-fixed point $c$ in $M_{\mathbb{R}}$. Since $1$ is not an eigenvalue of $g$, $c$ is unique. Suppose that a prime $p$ divides $\operatorname{ord}(g)$. Then $g^{\prime}=g^{\frac{\operatorname{ord}(g)}{p}}$ has order $p$ and the eigenvalues of the corresponding linear action of $g^{\prime}$ on $M_{\mathbb{C}}$ are primitive $p$th roots of unity. Let $c^{\prime}\in\frac{1}{p}M$ be a $g^{\prime}$-fixed point. Since $1$ is not an eigenvalue of $g^{\prime}$, $c^{\prime}$ is unique and hence $c=c^{\prime}$. If a prime $q\neq p$ also divides $\operatorname{ord}(g)$, then $c\in\frac{1}{p}M\cap\frac{1}{q}M=M$. ∎
@@
The complex representation ring $R(G)$ of $G$ is the free abelian group generated by isomorphism classes of finite-dimensional $\mathbb{C}G$-modules, modulo the relation $[V\oplus W]=[V]+[W]$ for $\mathbb{C}G$-modules $V,W$. Multiplication is defined by $[V][W]:=[V\otimes_{\mathbb{C}}W]$ and the identity element $1\in R(G)$ is the class $[\mathbf{1}]$ of the trivial representation. Let $\operatorname{Irr}(G)$ be the set of isomorphism classes of irreducible $\mathbb{C}G$-modules. Then $R(G)$ is a free $\mathbb{Z}$-module with basis $\{[V]:V\in\operatorname{Irr}(G)\}$. An element $r$ of $R(G)$ is effective if has the form $r=[V]$ for some $\mathbb{C}G$-module $V$. Equivalently, if we write $r=\sum_{V\in\operatorname{Irr}(G)}a_{V}[V]$ for some uniquely defined $a_{V}\in\mathbb{Z}$, then $r$ is effective if and only if $a_{V}\geq 0$ for all $V\in\operatorname{Irr}(G)$. We say that a power series in $R(G)[[t]]$ is effective if all its coefficients are effective. In computations, if $\chi$ is a complex representation of $G$, we will sometimes write $\chi\in R(G)$ instead of $[\chi]\in R(G)$ for convenience.
@@
Here $[\bigwedge^{m}V]=1\in R(G)$ when $m=0$.
@@
Given the affine representation $\rho$ above, the corresponding linear representation $\hat{\rho}$ (see Lemma 2.5) gives $M_{\mathbb{C}}$ the structure of a $\mathbb{C}G$-module, and, under the identification of $\operatorname{Aff}(M)$ as a subgroup of $\operatorname{GL}(\widetilde{M})$, $\rho$ gives $\widetilde{M}_{\mathbb{C}}$ the structure of a $\mathbb{C}G$-module. We have $\widetilde{M}_{\mathbb{C}}\cong M_{\mathbb{C}}\oplus\mathbf{1}$, $[\widetilde{M}_{\mathbb{C}}]=[M_{\mathbb{C}}]+1\in R(G)$ and $\det(I-\widetilde{M}_{\mathbb{C}}t)=(1-t)\det(I-M_{\mathbb{C}}t)\in R(G)[t]$.
@@
For any positive integer $m$, recall that the $m$th cyclotomic polynomial is the unique monic irreducible polynomial in $\mathbb{Z}[t]$ whose roots over $\mathbb{C}$ are precisely the primitive $m$th roots of unity. For each $g$ in $G$, the characteristic polynomial of $\rho(g)\in\operatorname{GL}(\widetilde{M})$ is defined over $\mathbb{Z}[t]$ and its roots over $\mathbb{C}$ are roots of unity of order $\operatorname{ord}(g)$. Hence the irreducible factors of the characteristic polynomial are cyclotomic polynomials.
@@ 2.4. Equivariant Ehrhart theory
Let $M\cong\mathbb{Z}^{d}$ be a lattice of rank $d$. Let $G$ be a finite group with identity element $\operatorname{id}$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation of $G$. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. We mention that if $P$ is a $G$-invariant lattice polytope of dimension less than $d$, one can reduce to the case above by replacing $M$ with its intersection with the affine span of $P$ (see Remark 2.12 below). Let $\widetilde{M}=M\oplus\mathbb{Z}$ and recall that we may view $\rho$ as a representation of $\widetilde{M}$ by identifying $\operatorname{Aff}(M)$ as the subgroup of $\operatorname{GL}(\widetilde{M})$ consisting of transformations that preserve projection $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ onto the last coordinate. Let $C_{P}\subset\widetilde{M}_{\mathbb{R}}$ denote the cone generated by $P\times\{1\}$ in $\widetilde{M}_{\mathbb{R}}$. Then $C_{P}$ is $G$-invariant. Recall from the introduction that for any positive integer $m$, $L(P,\rho;m)$ in $R(G)$ is the class of the permutation representation of $G$ acting on $P\cap\frac{1}{m}M$, or, equivalently, on $C_{P}\cap\widetilde{M}\cap\operatorname{pr}^{-1}(m)$. Then $L(P,\rho;m)$ is a quasi-polynomial of degree $d$ with coefficients in $R(G)\otimes_{\mathbb{Z}}\mathbb{Q}$ and constant term $1\in R(G)$ [Sta11a, Theorem 5.7]. For any $g$ in $G$, recall that we consider the rational polytope $P^{g}=\{x\in P:g\cdot x=x\}\subset M_{\mathbb{R}}$. Then the quasi-polynomial $L(P,\rho;m)(g)$ obtained by evaluating the associated character at $g$ in $G$ is equal to $L(P^{g};m)$ [Sta11a, Lemma 5.2]. That is, $L(P,\rho;m)$ precisely encodes the Ehrhart theory of the rational polytopes $\{P^{g}:g\in G\}$.
@@
We will also use the following equivalent formulation of equivariant Ehrhart theory. Let $W$ be a real affine space with corresponding real vector space $V$. Let $M^{\prime}$ be an affine lattice in $W$ with corresponding lattice $M\subset V$. Let $G$ be a finite group with an affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$. Let $P^{\prime}\subset W$ be a lattice polytope with respect to $M^{\prime}$ that is $G$-invariant. With this setup, we may define equivariant Ehrhart invariants as follows. Recall that we may identify $\operatorname{Aff}(M^{\prime})$ with $\{\phi\in\operatorname{Aff}(W):\phi(M^{\prime})=M^{\prime}\}$, and $\phi\in\operatorname{Aff}(W)$ gives rise to an $\mathbb{R}$-linear map $\bar{\phi}:V\to V$. Hence $V\otimes_{\mathbb{R}}\mathbb{C}$ has the structure of a $\mathbb{C}G$-module, which we denote $M_{\mathbb{C}}^{\prime}$. We let $\widetilde{M}_{\mathbb{C}}^{\prime}:=M_{\mathbb{C}}^{\prime}\oplus\mathbf{1}$ be the direct sum of $M_{\mathbb{C}}^{\prime}$ and the trivial $\mathbb{C}G$-module. For any positive integer $m$, define $L(P^{\prime},\rho^{\prime};m)$ in $R(G)$ to be the permutation representation of $G$ acting on $P^{\prime}\cap\frac{1}{m}M^{\prime}$. Let $\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)=1+\sum_{m>0}L(P^{\prime},\rho^{\prime};m)t^{m}\in R(G)[[t]]$ and $h^{*}(P^{\prime},\rho^{\prime};t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)\det(I-\widetilde{M}_{\mathbb{C}}^{\prime}t)\in R(G)[[t]]$. We claim that this formulation is equivalent to the formulation in the paper. Fix an element $u^{\prime}$ in $M^{\prime}$. Recall from Section 2.2 that we have a bijection $f_{u^{\prime}}:W\to M_{\mathbb{R}}$ such that $f_{u^{\prime}}(\frac{1}{m}M^{\prime})=\frac{1}{m}M$ for all positive integers $m$, and $f_{u^{\prime}}$ induces an isomorphism $\operatorname{Aff}(M^{\prime})\cong\operatorname{Aff}(M)$. Let $\rho:G\to\operatorname{Aff}(M)$ be the induced affine representation, and let $P=f_{u^{\prime}}(P^{\prime})$. Then $P$ is a lattice polytope with respect to $M$ that is $G$-invariant. By Example 2.11, $L(P,\rho;0)=1\in R(G)$ and, by Example 2.8, $\det(I-\widetilde{M}_{\mathbb{C}}t)=(1-t)\det(I-M_{\mathbb{C}}t)$. It then follows that $L(P,\rho;m)=L(P^{\prime},\rho^{\prime};m)$ for all positive integers $m$, $\operatorname{Ehr}(P,\rho;t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$ and $h^{*}(P,\rho;t)=h^{*}(P^{\prime},\rho^{\prime};t)$. Conversely, given an affine representation $\rho:G\to\operatorname{Aff}(M)$ and a $G$-invariant lattice polytope $P\subset M_{\mathbb{R}}$, we may regard $W:=M_{\mathbb{R}}$ as an real affine space with corresponding real vector space $V:=M_{\mathbb{R}}$, and $M^{\prime}:=M\subset W$ as an affine lattice in $W$ with corresponding lattice $M\subset V$. Let $\rho^{\prime}=\rho$ and $P^{\prime}=P$. Let $u^{\prime}$ in $M^{\prime}$ be the origin in $M$. Then $f_{u^{\prime}}:W\to M_{\mathbb{R}}$ is the identity map, $P=f_{u^{\prime}}(P^{\prime})$ and $f_{u^{\prime}}$ induces the equality $\operatorname{Aff}(M^{\prime})=\operatorname{Aff}(M)$. In particular, as above, $L(P,\rho;m)=L(P^{\prime},\rho^{\prime};m)$ for all positive integers $m$, $\operatorname{Ehr}(P,\rho;t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$ and $h^{*}(P,\rho;t)=h^{*}(P^{\prime},\rho^{\prime};t)$. Finally, in Section 3 and Section 4 we will construct lattice triangulations and consider commutative algebra invariants respectively associated to $P$, $\rho$ and $M$. Given an affine lattice $M^{\prime}$ together with an affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$ and $G$-invariant lattice polytope $P^{\prime}$, then, as above, one may choose a point $u^{\prime}$ in $M^{\prime}$ and consider the corresponding lattice structure on $M^{\prime}$, and hence consider the corresponding lattice triangulations and commutative algebra invariants. The choice of $u^{\prime}$ will not be important e.g. different choices of $u^{\prime}$ give isomorphic commutative algebra invariants. In particular, conditions (1), (2) and (3) in Theorem 1.4 are independent of the choice of $u^{\prime}$. For example, the semigroup $C_{P}\cap\widetilde{M}$ from the introduction is isomorphic to the semigroup $(P\cap(\cup_{m>0}\frac{1}{m}M^{\prime}))^{\times}$ considered in Section 2.2, and hence the commutative ring $S_{P}$ from the introduction is isomorphic to the semigroup algebra of $(P\cap(\cup_{m>0}\frac{1}{m}M^{\prime}))^{\times}$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Then the equivariant Ehrhart series $\operatorname{Ehr}(P,\rho;t)$ precisely encodes the equivariant $h^{*}$-series $h^{*}(P,\rho;t)$ together with the complex representation $\widetilde{M}_{\mathbb{C}}$ of $G$.
@@
The complex representation $\widetilde{M}_{\mathbb{C}}$ is determined by the eigenvalues of the action of the elements of $G$. Consider an element $g$ in $G$ acting on $\widetilde{M}_{\mathbb{C}}$ with eigenvalues $\{\lambda_{1},\ldots,\lambda_{d+1}\}$, where $d=\dim M_{\mathbb{R}}$. For any positive integer $m$, let $\operatorname{mult}(g,m):=|\{\lambda_{i}:\lambda_{i}^{m}=1\}|$ and $\operatorname{mult}^{\prime}(g,m):=|\{\lambda_{i}:\lambda_{i}^{m}=1,\lambda_{i}^{m^{\prime}}\neq 1\textrm{ for }m^{\prime}<m\}|$. Then $\operatorname{mult}(g,m)=\sum_{m^{\prime}|m}\operatorname{mult}^{\prime}(g,m^{\prime})$. By Remark 2.9, the multiplicity of a primitive $m$th root of unity $\zeta_{m}$ as an eigenvalue of $\rho(g)\in\operatorname{GL}(\widetilde{M})$ is independent of the choice of $\zeta_{m}$ in $\mathbb{C}$. It follows that $\widetilde{M}_{\mathbb{C}}$ is determined by the numbers $\{\operatorname{mult}^{\prime}(g,m):g\in G,m\in\mathbb{Z}_{>0}\}$, and hence by the numbers $\{\operatorname{mult}(g,m):g\in G,m\in\mathbb{Z}_{>0}\}$, for example, via Möbius inversion . Since $g^{m}$ acts on $\widetilde{M}_{\mathbb{C}}$ with eigenvalues $\{\lambda_{1}^{m},\ldots,\lambda_{d+1}^{m}\}$, we have $\operatorname{mult}(g,m)=\operatorname{mult}(g^{m},1)$. Hence $\widetilde{M}_{\mathbb{C}}$ is determined by the numbers $\{\operatorname{mult}(g,1):g\in G\}$. By [Sta11a, Lemma 5.3], $\operatorname{mult}(g,1)$ equals $\dim P^{g}$. On the other hand, by Theorem 2.1, the order of $t=1$ as a pole of $\operatorname{Ehr}(P,\rho;t)(g)=\operatorname{Ehr}(P^{g};t)$ is equal to $\dim P^{g}+1$. ∎
@@
Suppose that $G=\mathbb{Z}/p\mathbb{Z}$ for a prime $p$ and that $G$ acts on $M$ with a unique $G$-fixed point $c$. For example, when $p=2$, $P$ is called centrally symmetric. We will show that $h^{*}(P,\rho;t)$ is effective and, moreover, its coefficients are permutation representations. This generalizes the case when $p=2$ and $c\in M$ [Sta11a, Section 11]. Let $\deg(c)$ be the degree of $M$ as a sublattice of $M+\mathbb{Z}c\subset M_{\mathbb{Q}}$. By Lemma 2.6, $c\in\frac{1}{p}M$, and hence
@@
Fix $g\neq\operatorname{id}$ in $G$. Then $\operatorname{Ehr}(P,\rho;t)(g)=\frac{1}{1-t^{\deg(c)}}.$ On the other hand, $g$ acts on $M_{\mathbb{C}}$ with eigenvalues equal to primitive $m$th roots of unity. By Remark 2.9, $\det(I-\widetilde{M}_{\mathbb{C}}t)(g)$ equals $(1-t)(1+t+\cdots+t^{p-1})^{\frac{d}{p-1}}$. In particular, $p-1$ divides $d$. Recall that $\chi_{\operatorname{reg}}$ denotes the regular representation of $G$, and both the regular representation and trivial representation are permutation representations. We deduce that
@@
It follows that the coefficients of $h^{*}(P;t)-(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$ are divisible by $p$, and $h^{*}(P,\rho;t)$ is effective if and only if $h^{*}(P;t)\geq(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$, where the inequality is taken coefficientwise. It remains to deduce the inequality. Fix a generator $\sigma$ of $G$. Let $\{u_{1},\ldots,u_{r}\}\in M$ be a minimal set of vertices of $P$ with the property that the orbits $G\cdot u_{1},\ldots,G\cdot u_{r}$ affinely span $M_{\mathbb{Q}}$. Let $V=M_{\mathbb{Q}}-c$ considered as a $\mathbb{Q}G$-module. Let $V_{i}\subset V$ be the linear span of $(G\cdot u_{i})-c$. We have an isomorphism of $\mathbb{Q}G$-modules, $\phi_{i}:V_{i}\to\mathbb{Q}^{p}/\mathbb{Q}(1,\dots,1)$, $\phi_{i}(\sigma^{k}u_{i}-c)=e_{k}$, where $\sigma$ acts on the right hand side by cyclically shifting coordinates. The characteristic polynomial of $\sigma$ acting on $V_{i}$ is $1+t+\cdots+t^{p-1}$, an irreducible polynomial in $\mathbb{Q}[t]$, and hence $V_{i}$ is an irreducible $\mathbb{Q}G$-module. Minimality then implies that $V_{i}\cap V_{j}=\{0\}$ for all $i\neq j$, and $V=V_{1}\oplus\cdots\oplus V_{r}$. Let $Q\subset P$ be the convex hull of $G\cdot u_{1},\ldots,G\cdot u_{r}$ in $M_{\mathbb{R}}$. Stanley’s monotonicity theorem (the nonequivariant case of Theorem 1.7) implies that $h^{*}(Q;t)\leq h^{*}(P;t)$ [Sta93, Theorem 3.3]. Let $M^{\prime}$ be the affine sublattice of $M$ spanned by $G\cdot u_{1},\ldots,G\cdot u_{r}$, together with $c$ in the case when $c\in M$. Explicitly,
@@
Let $Q^{\prime}$ be $Q$ considered as a lattice polytope with respect to $M^{\prime}\subset M$. Then $h^{*}(Q^{\prime};t)\leq h^{*}(Q;t)$ by lattice-monotonicity of lattice polytopes [BKN23, Corollary 2.21]. We may directly compute that $h^{*}(Q^{\prime};t)=(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$ using [BJM13, Corollary 5.9], and the inequality follows. We mention that when $c\in M$, the inequality $h^{*}(P;t)\geq(1+t+\cdots+t^{p-1})^{\frac{d}{p-1}}$ may also be deduced from [Kar08, Corollary 1.2] and [AN99, Theorem 1.2]. Furthermore, when $c\in M$ and $p=2$, this inequality is originally due to [BHW07, Remark 1.6].

### v3

@@ 2.2. Affine transformations
[…]
@@ 2.3. Representation theory of finite groups
By Lemma 2.5, there is a $g$-fixed point $c$ in $M_{\mathbb{R}}$. Since $1$ is not an eigenvalue of $g$, $c$ is unique. Suppose that a prime $p$ divides $\operatorname{ord}(g)$. Then $g^{\prime}=g^{\frac{\operatorname{ord}(g)}{p}}$ has order $p$ and the eigenvalues of the corresponding linear action of $g^{\prime}$ on $M_{\mathbb{C}}$ are primitive $p$th roots of unity. Let $c^{\prime}\in\frac{1}{p}M$ be a $g^{\prime}$-fixed point. Since $1$ is not an eigenvalue of $g^{\prime}$, $c^{\prime}$ is unique and hence $c=c^{\prime}$. If a prime $q\neq p$ also divides $\operatorname{ord}(g)$, then $c\in\frac{1}{p}M\cap\frac{1}{q}M=M$.
@@
The complex representation ring $R(G)$ of $G$ is the free abelian group generated by isomorphism classes of finite-dimensional $\mathbb{C}G$-modules, modulo the relation $V\oplus W]=[V]+[W]$ for $\mathbb{C}G$-modules $V,W$. Multiplication is defined by $V][W]:=[V\otimes_{\mathbb{C}}W]$ and the identity element $1\in R(G)$ is the class $\mathbf{1}]$ of the trivial representation. Let $\operatorname{Irr}(G)$ be the set of isomorphism classes of irreducible $\mathbb{C}G$-modules. Then $R(G)$ is a free $\mathbb{Z}$-module with basis $\{[V]:V\in\operatorname{Irr}(G)\}$. An element $r$ of $R(G)$ is effective if has the form $r=[V]$ for some $\mathbb{C}G$-module $V$. Equivalently, if we write $r=\sum_{V\in\operatorname{Irr}(G)}a_{V}[V]$ for some uniquely defined $a_{V}\in\mathbb{Z}$, then $r$ is effective if and only if $a_{V}\geq 0$ for all $V\in\operatorname{Irr}(G)$. We say that a power series in $R(G)[[t]]$ is effective if all its coefficients are effective. In computations, if $\chi$ is a complex representation of $G$, we will sometimes write $\chi\in R(G)$ instead of $\chi]\in R(G)$ for convenience.
@@
Here $\bigwedge^{m}V]=1\in R(G)$ when $m=0$.
@@
[…]
@@
[…]
@@ 2.4. Equivariant Ehrhart theory
[…]
@@
We will also use the following equivalent formulation of equivariant Ehrhart theory. Let $W$ be a real affine space with corresponding real vector space $V$. Let $M^{\prime}$ be an affine lattice in $W$ with corresponding lattice $M\subset V$. Let $G$ be a finite group with an affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$. Let $P^{\prime}\subset W$ be a lattice polytope with respect to $M^{\prime}$ that is $G$-invariant.
With this setup, we may define equivariant Ehrhart invariants as follows. Recall that we may identify $\operatorname{Aff}(M^{\prime})$ with $\{\phi\in\operatorname{Aff}(W):\phi(M^{\prime})=M^{\prime}\}$, and $\phi\in\operatorname{Aff}(W)$ gives rise to an $\mathbb{R}$-linear map $\bar{\phi}:V\to V$. Hence $V\otimes_{\mathbb{R}}\mathbb{C}$ has the structure of a $\mathbb{C}G$-module, which we denote $M_{\mathbb{C}}^{\prime}$. We let $\widetilde{M}_{\mathbb{C}}^{\prime}:=M_{\mathbb{C}}^{\prime}\oplus\mathbf{1}$ be the direct sum of $M_{\mathbb{C}}^{\prime}$ and the trivial $\mathbb{C}G$-module. For any positive integer $m$, define $L(P^{\prime},\rho^{\prime};m)$ in $R(G)$ to be the permutation representation of $G$ acting on $P^{\prime}\cap\frac{1}{m}M^{\prime}$. Let $\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)=1+\sum_{m>0}L(P^{\prime},\rho^{\prime};m)t^{m}\in R(G)[[t]]$ and $h^{*}(P^{\prime},\rho^{\prime};t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)\det(I-\widetilde{M}_{\mathbb{C}}^{\prime}t)\in R(G)[[t]]$.
We claim that this formulation is equivalent to the formulation in the paper. Fix an element $u^{\prime}$ in $M^{\prime}$. Recall from Section 2.2 that we have a bijection $f_{u^{\prime}}:W\to M_{\mathbb{R}}$ such that $f_{u^{\prime}}(\frac{1}{m}M^{\prime})=\frac{1}{m}M$ for all positive integers $m$, and $f_{u^{\prime}}$ induces an isomorphism $\operatorname{Aff}(M^{\prime})\cong\operatorname{Aff}(M)$. Let $\rho:G\to\operatorname{Aff}(M)$ be the induced affine representation, and let $P=f_{u^{\prime}}(P^{\prime})$. Then $P$ is a lattice polytope with respect to $M$ that is $G$-invariant. By Example 2.11, $L(P,\rho;0)=1\in R(G)$ and, by Example 2.8, $\det(I-\widetilde{M}_{\mathbb{C}}t)=(1-t)\det(I-M_{\mathbb{C}}t)$. It then follows that $L(P,\rho;m)=L(P^{\prime},\rho^{\prime};m)$ for all positive integers $m$, $\operatorname{Ehr}(P,\rho;t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$ and $h^{*}(P,\rho;t)=h^{*}(P^{\prime},\rho^{\prime};t)$.
Conversely, given an affine representation $\rho:G\to\operatorname{Aff}(M)$ and a $G$-invariant lattice polytope $P\subset M_{\mathbb{R}}$, we may regard $W:=M_{\mathbb{R}}$ as an real affine space with corresponding real vector space $V:=M_{\mathbb{R}}$, and $M^{\prime}:=M\subset W$ as an affine lattice in $W$ with corresponding lattice $M\subset V$. Let $\rho^{\prime}=\rho$ and $P^{\prime}=P$. Let $u^{\prime}$ in $M^{\prime}$ be the origin in $M$. Then $f_{u^{\prime}}:W\to M_{\mathbb{R}}$ is the identity map, $P=f_{u^{\prime}}(P^{\prime})$ and $f_{u^{\prime}}$ induces the equality $\operatorname{Aff}(M^{\prime})=\operatorname{Aff}(M)$. In particular, as above, $L(P,\rho;m)=L(P^{\prime},\rho^{\prime};m)$ for all positive integers $m$, $\operatorname{Ehr}(P,\rho;t)=\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$ and $h^{*}(P,\rho;t)=h^{*}(P^{\prime},\rho^{\prime};t)$.
@@
Finally, in Section 3 and Section 4 we will construct lattice triangulations and consider commutative algebra invariants respectively associated to $P$, $\rho$ and $M$. Given an affine lattice $M^{\prime}$ together with an affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$ and $G$-invariant lattice polytope $P^{\prime}$, then, as above, one may choose a point $u^{\prime}$ in $M^{\prime}$ and consider the corresponding lattice structure on $M^{\prime}$, and hence consider the corresponding lattice triangulations and commutative algebra invariants. The choice of $u^{\prime}$ will not be important e.g. different choices of $u^{\prime}$ give isomorphic commutative algebra invariants. In particular, conditions (1), (2) and (3) in Theorem 1.4 are independent of the choice of $u^{\prime}$. For example, the semigroup $C_{P}\cap\widetilde{M}$ from the introduction is isomorphic to the semigroup $(P\cap(\cup_{m>0}\frac{1}{m}M^{\prime}))^{\times}$ considered in Section 2.2, and hence the commutative ring $S_{P}$ from the introduction is isomorphic to the semigroup algebra of $(P\cap(\cup_{m>0}\frac{1}{m}M^{\prime}))^{\times}$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Then the equivariant Ehrhart series $\operatorname{Ehr}(P,\rho;t)$ precisely encodes the equivariant $h^{*}$-series $h^{*}(P,\rho;t)$ together with the complex representation $\widetilde{M}_{\mathbb{C}}$ of $G$.
@@
[…]
@@
Suppose that $G=\mathbb{Z}/p\mathbb{Z}$ for a prime $p$ and that $G$ acts on $M$ with a unique $G$-fixed point $c$. For example, when $p=2$, $P$ is called centrally symmetric. We will show that $h^{*}(P,\rho;t)$ is effective and, moreover, its coefficients are permutation representations. This generalizes the case when $p=2$ and $c\in M$ [Sta11a, Section 11].
Let $\deg(c)$ be the degree of $M$ as a sublattice of $M+\mathbb{Z}c\subset M_{\mathbb{Q}}$. By Lemma 2.6, $c\in\frac{1}{p}M$, and hence
@@
Fix $g\neq\operatorname{id}$ in $G$. Then $\operatorname{Ehr}(P,\rho;t)(g)=\frac{1}{1-t^{\deg(c)}}.$ On the other hand, $g$ acts on $M_{\mathbb{C}}$ with eigenvalues equal to primitive $m$th roots of unity. By Remark 2.9, $\det(I-\widetilde{M}_{\mathbb{C}}t)(g)$ equals $(1-t)(1+t+\cdots+t^{p-1})^{\frac{d}{p-1}}$. In particular, p1p-1 divides $d$. Recall that $\chi_{\operatorname{reg}}$ denotes the regular representation of $G$, and both the regular representation and trivial representation are permutation representations. We deduce that
@@
It follows that the coefficients of $h^{*}(P;t)-(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$ are divisible by $p$, and $h^{*}(P,\rho;t)$ is effective if and only if $h^{*}(P;t)\geq(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$, where the inequality is taken coefficientwise.
It remains to deduce the inequality. Fix a generator $\sigma$ of $G$. Let $\{u_{1},\ldots,u_{r}\}\in M$ be a minimal set of vertices of $P$ with the property that the orbits $G\cdot u_{1},\ldots,G\cdot u_{r}$ affinely span $M_{\mathbb{Q}}$. Let $V=M_{\mathbb{Q}}-c$ considered as a $\mathbb{Q}G$-module. Let $V_{i}\subset V$ be the linear span of $(G\cdot u_{i})-c$. We have an isomorphism of $\mathbb{Q}G$-modules, $\phi_{i}:V_{i}\to\mathbb{Q}^{p}/\mathbb{Q}(1,\dots,1)$, $\phi_{i}(\sigma^{k}u_{i}-c)=e_{k}$, where $\sigma$ acts on the right hand side by cyclically shifting coordinates. The characteristic polynomial of $\sigma$ acting on $V_{i}$ is $1+t+\cdots+t^{p-1}$, an irreducible polynomial in $\mathbb{Q}[t]$, and hence $V_{i}$ is an irreducible $\mathbb{Q}G$-module. Minimality then implies that $V_{i}\cap V_{j}=\{0\}$ for all $i\neq j$, and $V=V_{1}\oplus\cdots\oplus V_{r}$. Let $Q\subset P$ be the convex hull of $G\cdot u_{1},\ldots,G\cdot u_{r}$ in $M_{\mathbb{R}}$. Stanley’s monotonicity theorem (the nonequivariant case of Theorem 1.7) implies that $h^{*}(Q;t)\leq h^{*}(P;t)$ [Sta93, Theorem 3.3]. Let $M^{\prime}$ be the affine sublattice of $M$ spanned by $G\cdot u_{1},\ldots,G\cdot u_{r}$, together with $c$ in the case when $c\in M$. Explicitly,
@@
Let $Q^{\prime}$ be $Q$ considered as a lattice polytope with respect to $M^{\prime}\subset M$. Then $h^{*}(Q^{\prime};t)\leq h^{*}(Q;t)$ by lattice-monotonicity of lattice polytopes [BKN23, Corollary 2.21]. We may directly compute that $h^{*}(Q^{\prime};t)=(1+t+\cdots+t^{p-1})^{\frac{d+1-\deg(c)}{p-1}}$ using [BJM13, Corollary 5.9], and the inequality follows.
We mention that when $c\in M$, the inequality $h^{*}(P;t)\geq(1+t+\cdots+t^{p-1})^{\frac{d}{p-1}}$ may also be deduced from [Kar08, Corollary 1.2] and [AN99, Theorem 1.2]. Furthermore, when $c\in M$ and $p=2$, this inequality is originally due to [BHW07, Remark 1.6].